# Learning to Extract Rational Evidence via Reinforcement Learning for Retrieval-Augmented Generation

### v3

@@ Abstract
Retrieval-Augmented Generation (RAG) effectively improves the accuracy of Large Language Models (LLMs). However, retrieval noises significantly impact the quality of LLMs’ generation, necessitating the development of denoising mechanisms. Previous methods extract evidence straightforwardly without explicit thinking, which risks filtering out key clues and struggles with generalization. To this end, we propose LEAR, which learns to extract rational evidence by (1) explicitly reasoning to identify potential cues within retrieval contents first, and then (2) consciously extracting to avoid omitting any key cues helpful for answering questions. Specifically, we frame evidence reasoning and extraction into one unified response for end-to-end training; apply knowledge token masks for disentanglement to derive reasoning-based and extraction-based answers; and devise three types of verifiable reward functions, including answer, length, and format, to update the model via the policy optimization algorithm. Extensive experiments on three benchmark datasets show the effectiveness of LEAR, providing compact and high-quality evidence, improving the accuracy of downstream tasks, and promoting effective application in online RAG systems [^1] The code, data, and models will be available at https://github.com/HITsz-TMG/LEAR..
@@ 1 Introduction
In this work, we propose LEAR, which LEArns to extract Rational evidence via reinforcement learning for RAG. Specifically, LEAR frames evidence reasoning (enclosed within <reason> and </reason>) and extraction (enclosed within <extract> and </extract>) into one response for end-to-end training and generates multiple responses to collect trajectories that contain good and bad experiences from the policy. After that, LEAR applies two knowledge token masks on each response to disentangle evidence reasoning and extraction for deriving reasoning-based and extraction-based answers (enclosed within <answer> and </answer> tags) to assess their quality respectively. Finally, we devise three types of rule-based verifiable reward functions, including answer, length, and format, to guide model optimization via Group Relative Policy Optimization (GRPO) Shao et al. (2024). As such, LEAR extracts rational evidence by reasoning about retrieval contents and then consolidating key cues into a concise and useful narrative, which is fed into LLMs to generate more accurate output. Our main contributions can be summarized as follows:
@@
  • We propose a novel rational evidence extraction learning framework, LEAR, which learns to discern potential cues in retrieval contents and then consciously consolidates any key cues into a concise yet very helpful evidence.
  • We conduct extensive experiments on three benchmark datasets, where the results fully demonstrate the superiority of LEAR in terms of performance, generalization, efficiency, as well as robustness against retrieval noise.
@@ 2.2 Empirical Study
Figure 3: The overall system framework of the proposed LEAR and previous vanilla paradigm. (a) Compared to vanilla evidence, our rational evidence is more skilled in extracting key cues. (b) Existing works learn to extract vanilla evidence via supervised fine-tuning, where the output, i.e., oracle that contains key clues, is usually created by hand-crafted strategies, e.g., String Inclusion and Lexical Overlap in Wang et al. (2023). (c) The proposed framework LEAR incentivizes rational evidence extraction capability in the extractor $\pi_{\theta}$ via reinforcement learning.
@@ 3 Methodology
The overall framework of LEAR is illustrated in Figure 3. In this section, we first introduce the acquisition of rational evidence (§3.1). Then, we define three verifiable reward functions (§3.2). Finally, we describe the training of the extractor through online policy optimization in detail (§3.3).
@@ 3.2 Modeling Verifiable Reward
The reward function plays a key role in Reinforcement Learning (RL) Ouyang et al. (2022); Wang et al. (2024); DeepSeek-AI et al. (2025), which guides the optimization process towards desired properties. To train LEAR via RL, we design three types of rule-based verifiable reward functions w.r.t. three primary desired properties for evidence extraction: (1) The correctness of the generated outputs; (2) The comprehensiveness of rationale and the conciseness of rational evidence; and (3) The format specification of the responses. For reward modeling, we first collect a set of Preao-tuple $<P,r,e,a,o_{r,e,f}>$, where a Preao-tuple e is composed of retrieved Passgaes $P$, rationale $r$, rational evidence $e$, golden answer $a$, generated output $o_{r,e,f}$ conditioned on different contexts. Given that, we design three types of verifiable reward functions to assess the rational evidence extractor w.r.t. the three desired properties:
@@ 3.3 Online Policy Optimization
Having obtained the final rewards of each response, we use the policy optimization algorithm GRPO [^4] Here, we adopt GRPO instead of Proximal Policy Optimization (PPO) Schulman et al. (2017), because PPO needs an  additional  critic  model  to  evaluate  policy  performance. Shao et al. (2024); DeepSeek-AI et al. (2025) to optimize LEAR, incentivizing rational evidence extraction capability in it via reinforcement learning. Specifically, for each input question $q$, GRPO first samples a group of responses $\mathcal{Y}=\{y_{1},y_{2},...,y_{G}\}$, where $G$ is the group size and each response consists of the rationale $r$, rational evidence $e$, and three outputs $o_{r}$, $o_{e}$, and $o_{f}$3.1). Subsequently, GRPO evaluates these responses based on the verifiable reward functions (§3.2) and obtains final rewards for each response =$\mathcal{R}=\{R_{1},R_{2},...,R_{G}\}$, where we omit the superscript “${final}$” for brevity. Different from PPO Schulman et al. (2017), GRPO directly compares the final rewards of candidate responses within the same group without needing an additional critic model. The advantage of the $i$-th response is determined through normalizing its reward $R_{i}$ using the mean and the standard deviation of rewards $\mathcal{R}$:
@@ 4 Experiment
In this section, we conduct extensive experiments on three knowledge-intensive benchmark datasets to answer the following Research Questions (RQs): RQ1: How does rational evidence perform, w.r.t. downstream task performance, compared to other vanilla evidence? RQ2: How do the properties of rational evidence vary with the RL training process? RQ3: How does the inference efficiency of LEAR compare with that of different types of methods? RQ4: Can rational evidence perform robustly against retrieval noises? RQ5: How do different parts of the answer rewards contribute to the final model performance?
@@ 4.1 Experimental Setup
Zero Full SeleCtx LLMLingua-2 Recomp FilCo SEER CoT LEAR
@@
Comparison Baselines.  To verify the effectiveness of LEAR, we compare it with the following three groups of competitive baselines: (1) Without Refinement (WR) includes (i) Zero-shot (Zero) generates output relying on LLMs’ parametric knowledge; (ii) Full Passage (Full) feeds all retrieval passages into LLMs; (2) Vanilla Refinement (VAR) directly extracts evidence without explicit thinking, which includes (i) Select Context (SeleCtx) Li et al. (2023) identifies and prunes redundancy in the input context based on perplexity; (ii) LLMLingua-2 Pan et al. (2024) distills compression knowledge from GPT-4 to reduce crucial information losing; (iii) FilCo Wang et al. (2023) trains a context filtering model to identify key clues; (iv) BottleNeck Zhu et al. (2024) applies the information bottle theory to select SFT data used to optimize filtering; (v) SEER Zhao et al. (2024a) learns to extract desired evidence via self-aligned learning; (3) Rational Refinement (RAR) includes (i) Chain-of-Thought (CoT) Wei et al. (2022) generates query-related information from retrieval passages with explicit thinking.
@@ 4.2 Main Comparison (RQ1)
The overall comparison results on NQ, TQA, and HotpotQA are shown in Table 4. From the experimental results, we mainly have the following observations. (1) In all cases (18/18), LEAR achieves the best or second-best results, indicating the superiority of supplementing RAG with rational evidence. (2) It is surprising that directly employing LEAR trained on NQ and tested on OOD datasets (i.e., TQA and HotpotQA) yields such impressive performance. This demonstrates that online policy optimization endows LEAR with superior generalization. (3) Compared to ‘Full’, LEAR possesses an extremely high compression ratio (e.g., 38.1x with 1.5B size model on NQ) and its performance is very close to or even better than that of ‘Full’. Though the baseline models have relatively high compression ratios, they are accompanied by significant performance degradation. This demonstrates again that rational evidence benefits downstream performance than vanilla one and also improves the signal-to-noise ratio. (4) By comparing the 1.5B LEAR model with the 7B one, we find that the 7B one tends to extract more informative evidence than the 1.5B one. (5) LEAR considerably outperforms VAR methods in almost all cases and provides more compact evidence, demonstrating the necessity of explicit evidence reasoning. (6) In HotpotQA, RAR methods significantly outperform VAR ones, and even better than ‘Full’, indicating that rational refinement is important to multi-hop question answering. In conclusion, the above results and observations fully validate the effectiveness and efficiency of rational evidence.
@@ 4.4 Inference Efficiency (RQ3)
Table 5 presents the inference latency of LEAR compared with FilCo (a representative VAR method) and CoT (a representative RAR method) on the 1.5B model within a 1$\times$×A800 station, where the evaluation batch size and max new tokens are set as 64 and 768, respectively. It can be observed that the time required for evidence extraction using LEAR is considerably shorter than FilCo and CoT, facilitating effective application in online RAG systems. Additionally, it is surprising that the average inference latency of FilCo is slightly longer than that of CoT, while FilCo generates fewer new tokens on average. We find that the standard deviation in the lengths of the evidence generated by FilCo is significantly higher than that of COT. Specifically, the std of FilCo on NQ, TQA, and HotpotQA is 107.5, 136.1, and 81.7, while that of CoT is 67.9, 61.6, and 63.9. Due to the instability in the length of the generated evidence, FilCo takes longer in batching generation instead. In summary, the above results fully verify the efficiency of LEAR during deployment.
@@
Average 38.72 39.13 38.61 39.22 39.31 39.48
Table 6: Performance comparison (EM) of LEAR trained on different combinations of answer rewards.
@@ 4.5 Robustness Analysis (RQ4)
In real-world applications, RAG systems commonly suffer from data noise resulting from imperfect retrieval. To simulate this scenario, we randomly sample a certain number (i.e., 0, 2, 4, 6, and 8) of irrelevant passages for each test query, where each query is equipped with 5 retrieved relevant passages as well as sampled irrelevant passages. We experiment on the 1.5B model, and the experimental results are presented in Figure 6. The results show that adding noise considerably degrades the performance of FilCo, while the performance degradation of LEAR is relatively small, where the green line is always below the blue one. In particular, LEAR with 8 irrelevant psgs outperforms FilCo without noise. In conclusion, this fully verifies the robustness of LEAR against noise.
@@ 4.6 Ablation Study (RQ5)
To evaluate the impact of different answer rewards on model performance, we train 1.5B-sized LEAR models on different combinations of answer rewards with average weighting. As shown in Table 6, among models trained on a single answer reward, training with evidence answer reward $R^{ans}_{e}$ yields the highest average performance, demonstrating that the quality of evidence is the most important in optimizing rational evidence. For two-answer reward combinations, ‘Evidence+Full Ctx’ slightly outperforms ‘Evidence+Rationale’, with noticeable gains on NQ in particular, indicating that the optimization of ‘Full Ctx’ matters. Notably, incorporating all three answer rewards yields the best performance on NQ and HotpotQA, as well as the highest average score, highlighting the benefits of multi-faceted reward assessment. In conclusion, the above results clearly manifest the contribution and necessity of each answer reward.
@@ 6 Conclusion
In this paper, we first unveil the limitations of the vanilla evidence extraction paradigm and explore the potential of the rational evidence extraction paradigm to solve these limitations. Specifically, we propose LEAR, a rational evidence extraction reinforcement learning framework. In particular, we unify evidence reasoning and extraction into one unified response, and devise verifiable reward functions to guide the optimization of LEAR for compact and high-quality evidence, where knowledge token masking and std clipping are applied to avoid information leakage and stabilize RL training, respectively. Extensive experiments fully demonstrate the superiority of LEAR w.r.t. performance, generalization, and robustness.
@@
Despite our innovations and improvements, it does have limitations, especially cascaded generation between the rationale and the rational evidence. This indicates that the rationale must be generated before the rational evidence, which increases the inference latency to some extent. Although the results in Table 5 show that the inference efficiency of LEAR remains superior to that of the baseline, we believe that further improvements in reasoning efficiency are necessary to achieve even higher inference speed. For example, reasoning should be conducted only when necessary Jiang et al. (2025), i.e., generating rational evidence directly for easy instances. We leave it for future research.

### v4

@@ Abstract
Retrieval-Augmented Generation (RAG) effectively improves the accuracy of Large Language Models (LLMs). However, retrieval noises significantly impact the quality of LLMs’ generation, necessitating the development of denoising mechanisms. Previous methods extract evidence straightforwardly without explicit thinking, which risks filtering out key clues and struggles with generalization. To this end, we propose EviOmni, which learns to extract rational evidence by (1) explicitly reasoning to identify potential cues within retrieval contents first, and then (2) consciously extracting to avoid omitting any key cues helpful for answering questions. Specifically, we frame evidence reasoning and evidence extraction into one unified response for end-to-end training; apply knowledge token masks for disentanglement to derive reasoning-based and extraction-based answers; and devise three types of verifiable reward functions, including answer, length, and format, to update the model via the policy optimization algorithm. Extensive experiments on three benchmark datasets show the effectiveness of EviOmni, providing compact and high-quality evidence, improving the accuracy of downstream tasks, and promoting effective application in online RAG systems [^1] The code, data, and models will be available at https://github.com/HITsz-TMG/EviOmni..
@@ 1 Introduction
In this work, we propose EviOmni, which learns to extract rational evidence via reinforcement learning for RAG. Specifically, EviOmni frames evidence reasoning (enclosed within <reason> and </reason>) and extraction (enclosed within <extract> and </extract>) into one response for end-to-end training and generates multiple responses to collect trajectories that contain good and bad experiences from the policy. After that, EviOmni applies two knowledge token masks on each response to disentangle evidence reasoning and extraction for deriving reasoning-based and extraction-based answers (enclosed within <answer> and </answer> tags) to assess their quality respectively. Finally, we devise three types of rule-based verifiable reward functions, including answer, length, and format, to guide model optimization via Group Relative Policy Optimization (GRPO) Shao et al. (2024). As such, EviOmni extracts rational evidence by reasoning about retrieval contents and then consolidating key cues into a concise and useful narrative, which is fed into LLMs to generate more accurate output. Our main contributions can be summarized as follows:
@@
  • We propose a novel rational evidence extraction learning framework, EviOmni, learning to discern potential cues in retrieval contents and consciously consolidates any key cues into a concise yet very helpful evidence.
  • We conduct extensive experiments on three benchmark datasets, where the results fully demonstrate the superiority of EviOmni in terms of performance, generalization, efficiency, and robustness against retrieval noise.
@@ 2.2 Empirical Study
Figure 3: The overall system framework of the proposed EviOmni and previous vanilla paradigm. (a) Compared to vanilla evidence, our rational evidence is more skilled in extracting key cues. (b) Existing works learn to extract vanilla evidence via supervised fine-tuning, where the output, i.e., oracle that contains key clues, is usually created by hand-crafted strategies, e.g., String Inclusion and Lexical Overlap in Wang et al. (2023). (c) The proposed framework EviOmni incentivizes rational evidence extraction capability in the extractor $\pi_{\theta}$ through RL.
@@ 3 Methodology
The overall framework of EviOmni is illustrated in Figure 3. In this section, we first introduce the acquisition of rational evidence (§3.1). Then, we define three verifiable reward functions (§3.2). Finally, we describe the training of the extractor through online policy optimization in detail (§3.3).
@@ 3.2 Modeling Verifiable Reward
The reward function plays a key role in Reinforcement Learning (RL) Ouyang et al. (2022); Wang et al. (2024); DeepSeek-AI et al. (2025), which guides the optimization process towards desired properties. To train EviOmni via RL, we design three types of rule-based verifiable reward functions w.r.t. three primary desired properties for evidence extraction: (1) The correctness of the generated outputs; (2) The comprehensiveness of rationale and the conciseness of rational evidence; and (3) The format specification of the responses. For reward modeling, we first collect a set of Preao-tuple $<P,r,e,a,o_{r,e,f}>$, where a Preao-tuple e is composed of retrieved Passgaes $P$, rationale $r$, rational evidence $e$, golden answer $a$, generated output $o_{r,e,f}$ conditioned on different contexts. Given that, we design three types of verifiable reward functions to assess the rational evidence extractor w.r.t. the three desired properties:
@@ 3.3 Online Policy Optimization
Having obtained the final rewards of each response, we use the policy optimization algorithm GRPO [^4] Here, we adopt GRPO instead of Proximal Policy Optimization (PPO) Schulman et al. (2017), because PPO needs an  additional  critic  model  to  evaluate  policy  performance. Shao et al. (2024); DeepSeek-AI et al. (2025) to optimize EviOmni, incentivizing rational evidence extraction capability in it via reinforcement learning. Specifically, for each input question $q$, GRPO first samples a group of responses $\mathcal{Y}=\{y_{1},y_{2},...,y_{G}\}$, where $G$ is the group size and each response consists of the rationale $r$, rational evidence $e$, and three outputs $o_{r}$, $o_{e}$, and $o_{f}$3.1). Subsequently, GRPO evaluates these responses based on the verifiable reward functions (§3.2) and obtains final rewards for each response =$\mathcal{R}=\{R_{1},R_{2},...,R_{G}\}$, where we omit the superscript “${final}$” for brevity. Different from PPO Schulman et al. (2017), GRPO directly compares the final rewards of candidate responses within the same group without needing an additional critic model. The advantage of the $i$-th response is determined through normalizing its reward $R_{i}$ using the mean and the standard deviation of rewards $\mathcal{R}$:
@@ 4 Experiment
In this section, we conduct extensive experiments on three knowledge-intensive benchmark datasets to answer the following Research Questions (RQs): RQ1: How does rational evidence perform, w.r.t. downstream task performance, compared to other vanilla evidence? RQ2: How do the properties of rational evidence vary with the RL training process? RQ3: How does the inference efficiency of EviOmni compare with that of different types of methods? RQ4: Can rational evidence perform robustly against retrieval noises? RQ5: How do different parts of the answer rewards contribute to the final model performance?
@@ 4.1 Experimental Setup
Zero Full SeleCtx LLMLingua-2 Recomp FilCo SEER CoT EviOmni
@@
Comparison Baselines.  To verify the effectiveness of EviOmni, we compare it with the following three groups of competitive baselines: (1) Without Refinement (WR) includes (i) Zero-shot (Zero) generates output relying on LLMs’ parametric knowledge; (ii) Full Passage (Full) feeds all retrieval passages into LLMs; (2) Vanilla Refinement (VAR) directly extracts evidence without explicit thinking, which includes (i) Select Context (SeleCtx) Li et al. (2023) identifies and prunes redundancy in the input context based on perplexity; (ii) LLMLingua-2 Pan et al. (2024) distills compression knowledge from GPT-4 to reduce crucial information losing; (iii) FilCo Wang et al. (2023) trains a context filtering model to identify key clues; (iv) BottleNeck Zhu et al. (2024) applies the information bottle theory to select SFT data used to optimize filtering; (v) SEER Zhao et al. (2024a) learns to extract desired evidence via self-aligned learning; (3) Rational Refinement (RAR) includes (i) Chain-of-Thought (CoT) Wei et al. (2022) generates query-related information from retrieval passages with explicit thinking.
@@ 4.2 Main Comparison (RQ1)
The overall comparison results on NQ, TQA, and HotpotQA are shown in Table 4. From the experimental results, we mainly have the following observations. (1) In all cases (18/18), EviOmni achieves the best or second-best results, indicating the superiority of supplementing RAG with rational evidence. (2) It is surprising that directly employing EviOmni trained on NQ and tested on OOD datasets (i.e., TQA and HotpotQA) yields such impressive performance. This demonstrates that online policy optimization endows EviOmni with superior generalization. (3) Compared to ‘Full’, EviOmni possesses an extremely high compression ratio (e.g., 38.1x with 1.5B size model on NQ) and its performance is very close to or even better than that of ‘Full’. Though the baseline models have relatively high compression ratios, they are accompanied by significant performance degradation. This demonstrates again that rational evidence benefits downstream performance than vanilla one and also improves the signal-to-noise ratio. (4) By comparing the 1.5B EviOmni model with the 7B one, we find that the 7B one tends to extract more informative evidence than the 1.5B one. (5) EviOmni considerably outperforms VAR methods in almost all cases and provides more compact evidence, demonstrating the necessity of explicit evidence reasoning. (6) In HotpotQA, RAR methods significantly outperform VAR ones, and even better than ‘Full’, indicating that rational refinement is important to multi-hop question answering. In conclusion, the above results and observations fully validate the effectiveness and efficiency of rational evidence.
@@ 4.4 Inference Efficiency (RQ3)
Table 5 presents the inference latency of EviOmni compared with FilCo (a representative VAR method) and CoT (a representative RAR method) on the 1.5B model within a 1$\times$×A800 station, where the evaluation batch size and max new tokens are set as 64 and 768, respectively. It can be observed that the time required for evidence extraction using EviOmni is considerably shorter than FilCo and CoT, facilitating effective application in online RAG systems. It is surprising that the average inference latency of FilCo is slightly longer than that of CoT, while FilCo generates fewer new tokens on average. We find that the standard deviation in the lengths of the evidence generated by FilCo is significantly higher than that of COT. Specifically, the std of FilCo on NQ, TQA, and HotpotQA is 107.5, 136.1, and 81.7, while that of CoT is 67.9, 61.6, and 63.9. Due to the instability in the length of the generated evidence, FilCo takes longer in batching generation instead. In all, the above results fully verify the efficiency of EviOmni during deployment.
@@
Average 39.05 39.13 38.61 39.22 39.31 39.48
Table 6: Performance comparison (EM) of EviOmni trained on different combinations of answer rewards.
@@ 4.5 Robustness Analysis (RQ4)
In real-world applications, RAG systems commonly suffer from data noise resulting from imperfect retrieval. To simulate this scenario, we randomly sample a certain number (i.e., 0, 2, 4, 6, and 8) of irrelevant passages for each test query, where each query is equipped with 5 retrieved relevant passages and sampled irrelevant passages. We experiment on the 1.5B model, and the experimental results are presented in Figure 6. The results show that adding noise considerably degrades the performance of FilCo, while the performance degradation of EviOmni is relatively small, where the green line is always below the blue one. In particular, EviOmni with 8 irrelevant psgs outperforms FilCo without noise. In all, this fully verifies the robustness of EviOmni to noise.
@@ 4.6 Ablation Study (RQ5)
To evaluate the impact of different answer rewards on model performance, we train 1.5B-sized EviOmni models on different combinations of answer rewards with average weighting. As shown in Table 6, among models trained on a single answer reward, training with evidence answer reward $R^{ans}_{e}$ yields the highest average performance, demonstrating that the quality of evidence is the most important in optimizing rational evidence. For two-answer reward combinations, ‘Evidence+Full Ctx’ slightly outperforms ‘Evidence+Rationale’, with significant gains on NQ, indicating that the optimization of ‘Full Ctx’ matters. Notably, incorporating all three answer rewards yields the best performance on NQ and HotpotQA, and the highest average score, indicating the benefits of multi-faceted reward assessment. In all, the above results manifest the contribution and necessity of each answer reward.
@@ 6 Conclusion
In this paper, we first unveil the limitations of the vanilla evidence extraction paradigm and explore the potential of the rational evidence extraction paradigm to solve these limitations. Specifically, we propose EviOmni, a rational evidence extraction reinforcement learning framework. In particular, we unify evidence reasoning and extraction into one unified response, and devise verifiable reward functions to guide the optimization of EviOmni for compact and high-quality evidence, where knowledge token masking and std clipping are applied to avoid information leakage and stabilize RL training, respectively. Extensive experiments fully manifest the superiority of EviOmni w.r.t. performance, generalization, and robustness.
@@
Despite our innovations and improvements, it does have limitations, especially cascaded generation between the rationale and the rational evidence. This indicates that the rationale must be generated before the rational evidence, which increases the inference latency to some extent. Although the results in Table 5 show that the inference efficiency of EviOmni remains superior to that of the baseline, we believe that further improvements in reasoning efficiency are necessary to achieve even higher inference speed. For example, reasoning should be conducted only when necessary Jiang et al. (2025), i.e., generating rational evidence directly for easy instances. We leave it for future research.